OpenStudy (anonymous):

Can someone refresh my mind on how powers work? Such as 10 to the -4th power

OpenStudy (lgbasallote):

\[\huge a^{-m} \implies \frac 1{a^m}\] does that help?

OpenStudy (lgbasallote):

\[\LARGE 10^{-4} \implies \frac 1{10^4} \implies \frac 1{10 \times 10 \times 10 \times 10}\]

OpenStudy (anonymous):

some can you finish the problem?

OpenStudy (lgbasallote):

this guy solves faster than me http://www.wolframalpha.com/input/?i=10%5E%28-4%29

OpenStudy (anonymous):

thank you